# MVG4D: Image Matrix-Based Multi-View and Motion Generation for 4D Content Creation from a Single Image

### v1

@@ 1 Introduction
With the rapid development of generative modeling technologies, the ability to create various digital contents such as 2D images [Rombach et al.(2022)], videos [Blattmann et al.(2023), Chai et al.(2023), Esser et al.(2023)], and 3D scenes [Long et al.(2024), Liu et al.(2024)] has been significantly improved. However, generating continuous and high-quality dynamic 4D scenes [Singer et al.(2023), Ling et al.(2024)] remains challenging. Dynamic 4D scenes are typically represented by dynamic Neural Radiance Fields (NeRF) [Mildenhall et al.(2021)] and aim to display consistent appearance, geometry, and motion from any viewpoint. Recent works have introduced time-set variables to optimize Gaussian deformation fields based on the 3D Gaussian Splatting (3D GS) [Kerbl et al.(2023)] technique, resulting in a novel 4D representation called 4D Gaussian Splatting (4D GS) [Wu et al.(2024)]. This method aims to achieve efficient training and storage for real-time dynamic scene rendering while maintaining high-quality output.
Dynamic 4D scene generation often relies on the concept of NeRF, which accurately reproduces the appearance, geometric structure, and motion state of objects from any angle. This technique combines the advantages of video processing and 3D modeling to generate high-quality content ranging from simple text and video descriptions to complex 4D scenes. For instance, Hexplane [Cao and Johnson(2023)] technology enables the generation of dynamic videos directly from text, which can then be converted into 4D presentations through advanced diffusion models, greatly enriching content creation possibilities. Emerging technologies such as CONSISTENT4D [Jiang et al.(2024)] and MAV3D [Singer et al.(2023)], attempt to generate coherent 4D scenes from a single video source by optimizing the NeRF framework and diffusion models. These methods effectively combine video information with dynamic 3D models to generate 4D content rich in details and dynamic textures. Additionally, techniques such as Animate124 [Zhao et al.(2023)] explore the transformation of a single image into a dynamic 3D video, further expanding the applications of dynamic scene generation.
@@
The 4D GS method employs a lightweight MLP to predict 3D Gaussian deformations for new timestamps, offering a novel display representation for dynamic scenes. This approach significantly reduces optimization time and enables real-time processing while maintaining high-quality deformable Gaussian Splatting.
@@ 2 Related Work
2.0.1 Single View to Multi-view Image Generation
Exciting progress has been made in using diffusion models to generate images from multiple views using a single image of an object. For example, Zero123 [Liu et al.(2023b)] utilizes the geometric prior learned by a large-scale diffusion model to train a conditional diffusion model on synthetic datasets, achieving zero-shot synthesis of new perspective images from single-view inputs. By combining images from different perspectives as target images, Zero123++ [Shi et al.(2023)] significantly overcomes the problem of no correlation between the images generated by Zero123, and improves the quality and consistency of generating consistent multi-perspective images from a single image.
@@ 2.0.1 Single View to Multi-view Image Generation
2.0.2 Image-to-3D Generation
The technology for generating 3D models from images has been rapidly maturing, and many 3D models representation methods based on Neural Radiance Fields (NeRF), point clouds, meshes, and Gaussian Splatting have emerged. Most of these methods, which are built from multi-view images, focus on creating a 3D representation of the scene. One-2-3-45++ [Liu et al.(2024)] combines a 2D diffusion model and a 3D native diffusion model for multi-view conditions to quickly generate 3D meshes through consistent multi-view image generation and 3D reconstruction. However, this traditional diffusion model-based scheme generally has the problem of long optimization time. 3D GS [Kerbl et al.(2023)] uses 3D Gaussian scene representation and real-time differentiable renderer to achieve real-time rendering of radiation field, which effectively reduces the time required for 3D content reconstruction and achieves high-quality scene representation by optimizing 3D Gaussian attributes and density control. While the task of real-time 3D reconstruction with just one image input has not yet been addressed, LRM [Hong et al.(2023)] uses a Transformer [Vaswani et al.(2017)]-based encoder-decoder architecture to predict Neural Radiance Fields (NeRF) directly from a single image for 3D reconstruction. This has significant implications for the task of 3D reconstruction based on a single image.
@@ 2.0.2 Image-to-3D Generation
2.0.3 4D Generation
Dynamic scene rendering, also known as 4D generation work, aims to enable real-time rendering of dynamic 3D scenes with efficient training and storage. In this field, neural radiance Fields (NeRF) techniques have been widely used to represent 4D scenes. For example, Neural-3D-Video [Li et al.(2022)] uses time-conditioned NeRF and a series of compact latent codes to represent dynamic scenes, and uses a hierarchical training scheme and light importance sampling to significantly improve the training speed and the perceptual quality of generated images. TiNeuVox [Fang et al.(2022)] accelerates the optimization of dynamic radiance fields by combining time-aware voxel features and micro-coordinate deformation networks to represent scenes. This work significantly reduces training time and storage costs while maintaining high-quality rendering. However, the robustness of dynamic radiance field reconstruction methods remains a challenge. RoDynRF’s [Liu et al.(2023a)] method for reconstructing camera trajectories and dynamic radiance fields from randomly captured dynamic monocular videos, enabling rendering that focuses on robust dynamic scenes. And MSTH [Wang et al.(2023)], which uses combinations of hash codes to represent dynamic scenes.
With the development of 3D Gaussian Splatting (3D GS), 4D Gaussian Splatting (4D GS) provides an innovative way to render dynamic scenes, which handles dynamic changes through continuous 4D representation. The 3D GS function is manipulated to adapt to temporal and spatial changes, thus enabling real-time rendering and high-resolution output. 4D GS emphasizes computation and storage efficiency, and utilizes a compact network structure to efficiently capture complex dynamic changes while maintaining high-quality rendering effects. Furthermore, DreamGaussian4D [Ren et al.(2023)] uses an image-to-3D framework to fit a static 3D GS function, and then optimalizes the dynamic representation by learning the motion that drives the video. Finally, the 4D GS is exported as an animated mesh sequence, and the texture map is optimized through a video-to-video process. Recently, Stable Video 4D (SV4D) [Xie et al.(2024)] provides an innovative approach for dynamic 3D content generation that enables 4D generation via multi-frame multi-view coherence. Unlike traditional approaches that rely on independently trained video generation and novel view synthesis models, SV4D employs a unified diffusion model that is able to generate multi-view videos of dynamic 3D objects from a mutil-view video.
@@ 4.1.1 Dataset
To evaluate the effectiveness of our method in generating high-quality dynamic 4D content, we conducted qualitative and qualitative evaluations based on datasets such as Objaverse [Deitke et al.(2023)].
@@ 4.1.2 Evaluation Metrics
For the task of generating 4D content from a single input image, designing appropriate evaluation metrics is essential to comprehensively assess the quality, consistency, and realism of the generated multi-view and dynamic content. Our method is evaluated using the following metrics: CLIP-I, PSNR, Time, FVD [Unterthiner et al.(2018)], each targeting different aspects of performance:
@@ 4.1.3 Baselines
To evaluate the quality of dynamic 4D content generation from a single image, we compare our method MVG4D with several recent 4D content generation methods, including DreamGaussian4D [Ren et al.(2023)], V4D [Gan et al.(2023)], 4Diffusion [Zhang et al.(2024)], etc. We use official code published by the respective authors to generate comparison results.
@@ 4.2 Implementation Details
RealFusion-V [Melas-Kyriazi et al.(2023)] 0.803
Animate124 [Zhao et al.(2023)] 0.854
DreamGaussian4D [Ren et al.(2023)] 0.923
Consistent4D [Jiang et al.(2024)] 0.921
EG4D [Sun et al.(2024)] 0.954
@@ 4.3.1 Image Similarity Metrics
TiNeuVox-B [Fang et al.(2022)] 32.67
DreamGaussian4D [Ren et al.(2023)] 34.05
KPlanes [Fridovich-Keil et al.(2023)] 31.61
V4D [Gan et al.(2023)] 33.72
4Diffusion [Zhang et al.(2024)] 35.07
@@
Consistent4D [Jiang et al.(2024)] 1133.93 741.52 871.95
SV3D [Voleti et al.(2025)] 989.53 526.78 690.49
STAG4D [Zeng et al.(2025)] 861.88 636.83 546.56
SV4D [Xie et al.(2024)] 677.68 525.65 614.35
Dreamgaussian4D [Ren et al.(2023)] 697.8 615.68 638.15
@@ 4.3.2 Time Efficiency Analysis
Method Time $\downarrow$
V4D [Gan et al.(2023)] 6.9 h
TiNeuVox-B [Fang et al.(2022)] 28 mins
DreamGaussian4D [Ren et al.(2023)] 13 mins
Stag4D [Zeng et al.(2025)] 9 mins
MVG4D (ours) 8 m 46 s
@@ 4.4 Qualitative Results
Accurate detail rendering and boundary handling are crucial for achieving high-quality dynamic 4D content generation. As shown in Figure 5, we compare the performance of the benchmark Stable Video Diffusion (SVD) [Blattmann et al.(2023)] method and our proposed MVG4D method. The SVD method exhibits noticeable blurring in high-frequency details, with significant noise and blurred transitions along edges, reducing clarity. In contrast, MVG4D achieves sharper boundaries by preserving structural details and effectively suppressing noise. For boundary handling, the SVD method produces a sawtooth effect along edges, resulting in uneven contours and abrupt transitions, whereas MVG4D delivers smoother and more natural results, enhancing overall visual fidelity.
@@
4.5 Ablation Study
4.5.1 Ablation Experiments on Viewpoint Adjustment

### v2

@@ 1 Introduction
With the rapid development of generative modeling technologies, the ability to create various digital contents such as 2D images [26], videos [1, 3, 7], and 3D scenes [21, 17, 4, 11] has been significantly improved. However, generating continuous and high-quality dynamic 4D scenes [28, 16] remains challenging. Dynamic 4D scenes are typically represented by Dynamic Neural Radiance Fields(NeRF) [23],which aim to model appearance, geometry, and motion across different viewpoints. For instance, Hexplane [2] technology enables the generation of dynamic videos directly from text, which can then be converted into 4D presentations through advanced diffusion models, greatly enriching content creation possibilities. Emerging technologies such as CONSISTENT4D [13] and MAV3D [28], attempt to generate coherent 4D scenes from a single video source by optimizing the NeRF framework and diffusion models. These methods effectively combine video information with dynamic 3D models to generate 4D content rich in details and dynamic textures. Additionally, techniques such as Animate124 [38] explore the transformation of a single image into a dynamic 3D video, further expanding the applications of dynamic scene generation.
@@
Recent works have introduced time-set variables to optimize Gaussian deformation fields based on the 3D Gaussian Splatting (3D GS) [14] technique, resulting in a novel 4D representation called 4D Gaussian Splatting (4D GS) [34]. This method aims to achieve efficient training and storage for real-time dynamic scene rendering while maintaining high-quality output. The 4D GS method employs a lightweight MLP to predict 3D Gaussian deformations for new timestamps, offering a novel display representation for dynamic scenes. This approach significantly reduces optimization time and enables real-time processing while maintaining high-quality deformable Gaussian Splatting.
@@ 2 Related Work
2.1 Single View to Multi-view Image Generation
Exciting progress has been made in using diffusion models to generate images from multiple views using a single image of an object. For example, Zero123 [18] utilizes the geometric prior learned by a large-scale diffusion model to train a conditional diffusion model on synthetic datasets, achieving zero-shot synthesis of new perspective images from single-view inputs. By combining images from different perspectives as target images, Zero123++ [27] significantly overcomes the problem of no correlation between the images generated by Zero123, and improves the quality and consistency of generating consistent multi-perspective images from a single image. SyncDreamer [20] further optimizes its adaptability to downstream 3D generation tasks on the basis of the previous two, but it still has shortcomings due to the high hardware requirements.
@@ 2.1 Single View to Multi-view Image Generation
2.2 Image-to-3D Generation
The technology for generating 3D models from images has been rapidly maturing, and many 3D models representation methods based on point clouds [4, 11], meshes, implicit neural representations [24, 5] and Gaussian Splatting have emerged. Most of these methods, which are built from multi-view images, focus on creating a 3D representation of the scene. One-2-3-45++ [17] combines a 2D diffusion model and a 3D native diffusion model for multi-view conditions to quickly generate 3D meshes through consistent multi-view image generation and 3D reconstruction. However, this traditional diffusion model-based scheme generally has the problem of long optimization time. 3D GS [14] uses 3D Gaussian scene representation and real-time differentiable renderer to achieve real-time rendering of radiation field, which effectively reduces the time required for 3D content reconstruction and achieves high-quality scene representation by optimizing 3D Gaussian attributes and density control. While the task of real-time 3D reconstruction with just one image input has not yet been addressed, LRM [12] uses a Transformer [31]-based encoder-decoder architecture to predict Neural Radiance Fields (NeRF) directly from a single image for 3D reconstruction. This has significant implications for the task of 3D reconstruction based on a single image.
@@ 2.2 Image-to-3D Generation
2.3 4D Generation
Dynamic scene rendering, also known as 4D generation work, aims to enable real-time rendering of dynamic 3D scenes with efficient training and storage. In this field, neural radiance Fields (NeRF) techniques have been widely used to represent 4D scenes. For example, Neural-3D-Video [15] uses time-conditioned NeRF and a series of compact latent codes to represent dynamic scenes, and uses a hierarchical training scheme and light importance sampling to significantly improve the training speed and the perceptual quality of generated images. TiNeuVox [8] accelerates the optimization of dynamic radiance fields by combining time-aware voxel features and micro-coordinate deformation networks to represent scenes. This work significantly reduces training time and storage costs while maintaining high-quality rendering. However, the robustness of dynamic radiance field reconstruction methods remains a challenge. RoDynRF’s [19] method for reconstructing camera trajectories and dynamic radiance fields from randomly captured dynamic monocular videos, enabling rendering that focuses on robust dynamic scenes. And MSTH [33], which uses combinations of hash codes to represent dynamic scenes.
With the development of 3D Gaussian Splatting (3D GS), 4D Gaussian Splatting (4D GS) provides an innovative way to render dynamic scenes, which handles dynamic changes through continuous 4D representation. The 3D GS function is manipulated to adapt to temporal and spatial changes, thus enabling real-time rendering and high-resolution output. 4D GS emphasizes computation and storage efficiency, and utilizes a compact network structure to efficiently capture complex dynamic changes while maintaining high-quality rendering effects. Furthermore, DreamGaussian4D [25] uses an image-to-3D framework to fit a static 3D GS function, and then optimalizes the dynamic representation by learning the motion that drives the video. Finally, the 4D GS is exported as an animated mesh sequence, and the texture map is optimized through a video-to-video process. Recently, Stable Video 4D (SV4D) [35] provides an innovative approach for dynamic 3D content generation that enables 4D generation via multi-frame multi-view coherence. Unlike traditional approaches that rely on independently trained video generation and novel view synthesis models, SV4D employs a unified diffusion model that is able to generate multi-view videos of dynamic 3D objects from a mutil-view video.
@@ 4.1.1 Dataset
To evaluate the effectiveness of our method in generating high-quality dynamic 4D content, we conducted qualitative and qualitative evaluations based on datasets such as Objaverse [6].
@@ 4.1.2 Evaluation Metrics
For the task of generating 4D content from a single input image, designing appropriate evaluation metrics is essential to comprehensively assess the quality, consistency, and realism of the generated multi-view and dynamic content. Our method is evaluated using the following metrics: CLIP-I, PSNR, Time, FVD [30], each targeting different aspects of performance:
@@ 4.1.3 Baselines
To evaluate the quality of dynamic 4D content generation from a single image, we compare our method MVG4D with several recent 4D content generation methods, including DreamGaussian4D [25], V4D [10], 4Diffusion [37], etc. We use official code published by the respective authors to generate comparison results.
@@ 4.2 Implementation Details
RealFusion-V [22] 0.803
Animate124 [38] 0.854
DreamGaussian4D [25] 0.923
Consistent4D [13] 0.921
EG4D [29] 0.954
@@ 4.3.1 Image Similarity Metrics
TiNeuVox-B [8] 32.67
DreamGaussian4D [25] 34.05
KPlanes [9] 31.61
V4D [10] 33.72
4Diffusion [37] 35.07
@@
Consistent4D [13] 1133.93 741.52 871.95
SV3D [32] 989.53 526.78 690.49
STAG4D [36] 861.88 636.83 546.56
SV4D [35] 677.68 525.65 614.35
Dreamgaussian4D [25] 697.8 615.68 638.15
@@
Method Time $\downarrow$
V4D [10] 6.9 h
TiNeuVox-B [8] 28 mins
DreamGaussian4D [25] 13 mins
Stag4D [36] 9 mins
MVG4D (ours) 8 m 46 s
@@ 4.4 Qualitative Results
Accurate detail rendering and boundary handling are crucial for achieving high-quality dynamic 4D content generation. As shown in Figure 4, we compare the performance of the benchmark Stable Video Diffusion (SVD) [1] method and our proposed MVG4D method. The SVD method exhibits noticeable blurring in high-frequency details, with significant noise and blurred transitions along edges, reducing clarity. In contrast, MVG4D achieves sharper boundaries by preserving structural details and effectively suppressing noise. For boundary handling, the SVD method produces a sawtooth effect along edges, resulting in uneven contours and abrupt transitions, whereas MVG4D delivers smoother and more natural results, enhancing overall visual fidelity.
4.5 Ablation Study
4.5.1 Ablation Experiments on Viewpoint Adjustment